OpenStudy (anonymous):

A 26-foot ladder is placed against a house and reaches the 24-foot roof. What is the distance between the base of the ladder and the base of the house if the ground meets the house at a right angle? 2 ft 4 ft 10 ft 8 ft

OpenStudy (anonymous):

wouldnt it be 26^2 + B^2 = 24^2?

OpenStudy (johnweldon1993):

 Not quite...since our hypotenuse will be the 26 that would be the 'c' in our equation \[\large a^2 + b^2 = c^2\] we have a and we have c \[\large 24^2 + b^2 = 26^2\]

OpenStudy (johnweldon1993):

So to solve for the 'b' we would do \[\large 26^2 - 24^2 = b^2\] \[\large b = \sqrt{26^2 - 24^2}\]

OpenStudy (anonymous):

(X²)=(26²)-(24²), so X= the square root of (676-576), so X= 10 ft ( choice C)

OpenStudy (anonymous):

answer is 10 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that is correct

OpenStudy (johnweldon1993):

As said above, correct! :)